OpenStudy (alexandervonhumboldt2):

I'm in a Sailing University in Russia and we were given a question which is a math question: I'm mot sure I did it correctly: A ship is going with speed v to the West. Known that the wind is blowing from south-west. The speed of the wind which is measured on the ship is u0. Find the speed of wind reletivly to Earth.

OpenStudy (alexandervonhumboldt2):

|vercot^a^|=|vector^v^|*sin45=(|vetor^v^|*sqrt2)/2. |vector^u0^|^2 - |vector^a^|^2=|vector^b^|^2. |vector^b^|^2=(|vector^u^|+|vector^v^|*cos45)^2=(|vector^u^|+(|vector^v^|*sqrt2)/2) |vector^u0^|^2 - |vector^a^|^2=(|vector^u^|+(|vector^v^|*sqrt2)/2) |vector^u^|+(|vector^v^|*sqrt2)/2=sqrt(vector^u0^|^2 - |vector^a^|^2)=sqrt(u0^2 - (V*sqrt2)/2) So the answer is \(\LARGE\color{darkgreen}{\text{u=sqrt(u0 - v^2/2) - V/sqrt2.}}\)

OpenStudy (alexandervonhumboldt2):

Am i Correct?

OpenStudy (alexandervonhumboldt2):

@rational

OpenStudy (alexandervonhumboldt2):

@perl

OpenStudy (alexandervonhumboldt2):

@Preetha

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (alexandervonhumboldt2):

what yes?

OpenStudy (alexandervonhumboldt2):

@texaschic101

OpenStudy (alexandervonhumboldt2):

did you meant i'm correct>?

OpenStudy (anonymous):

yeah sorry your correct

OpenStudy (alexandervonhumboldt2):

you sure? cause it is very importaint to me to pass this exam.

OpenStudy (alexandervonhumboldt2):

well yeah here is the diagamr i made

OpenStudy (alexandervonhumboldt2):

let me draw it

OpenStudy (alexandervonhumboldt2):

@perl

OpenStudy (perl):



OpenStudy (alexandervonhumboldt2):

yes that is 45

OpenStudy (alexandervonhumboldt2):

@perl here i finnaly posted my drawing seems we have different

OpenStudy (alexandervonhumboldt2):

lol got it.